# Hallmarks of Spin Textures for High-Harmonic Generation

### v1

@@ Abstract
Figure 1: Illustration of the patterns of the spin-texture in momentum space for the high-harmonics generation. Left panel: in 2D systems with spin texture invariant under twofold rotational symmetry, $\hat{\mathcal{C}}_{2}$, there is no emission of even-order harmonics. Right panel: for a spin texture exhibiting a $\hat{\mathcal{C}}_{2}$ symmetry breaking, the emission of even-order harmonics is allowed.
@@
This can be compared with the corresponding transformation properties of the spin texture, namely, $\boldsymbol{\sigma}_{{\bf k}_{\parallel},-{\bf k}_{\perp}}^{\parallel}=-\boldsymbol{\sigma}_{{\bf k}_{\parallel},{\bf k}_{\perp}}^{\parallel}$ and $\boldsymbol{\sigma}_{{\bf k}_{\parallel},-{\bf k}_{\perp}}^{\perp}=\boldsymbol{\sigma}_{{\bf k}_{\parallel},{\bf k}_{\perp}}^{\perp}$. From this point onward, we focus on even-order harmonics. Having identified the $\hat{\mathcal{C}}_{2}$-invariant spin texture as a sufficient condition for their suppression, we now examine scenarios where spin textures break this symmetry, enabling their emission. In this context, it is essential to distinguish between the two cases in which TR symmetry is either preserved or broken.
Figure 2: Spin textures and high-harmonic response for different physical configurations. Schematic illustration of the trigonal lattice (a) with threefold rotational symmetry and antiferromagnetic pattern (d) with broken twofold rotational symmetry. Both systems possess a vertical mirror symmetry $\mathcal{M}_{x}$ along the $x$-axis, and in both cases, the incident field (green arrow) is polarized perpendicular to this direction. Blue and orange arrows represent the expected polarizations for odd- and even-order harmonics, respectively. Panels (b) and (e) show the spin textures in momentum space corresponding to the physical configuration in (a) and (d), respectively. The color legend indicates the out-of-plane spin component $\langle\sigma_{z}\rangle$, with color blue, gray and red corresponding to $\langle\sigma_{z}\rangle=-1$, 00 and $1$, respectively. Black circles mark spots in the spin textures where $\hat{\mathcal{C}}_{2}$ symmetry, as described by Eq. (4), is broken. HHG spectra for the time-reversal invariant trigonal configuration (c) and for the antiferromagnetic pattern (f), respectively. The Hamiltonian parameters, in units of the hopping energy $t$, are: $\gamma_{R}=0.1$ for both cases; $\lambda=0.1$ in (c); and $\Delta\varepsilon=0.5$, $B=0.2$ in (f). The computations were performed at temperature $T=0$ with chemical potential $\mu=1.0$. In both cases, the incident field has intensity $|{\bf A}_{0}|=1.0$ and frequency $\Omega_{in}=0.5$. Polarizations of odd- (blue dots) and even-order (orange squares, with different shades of orange) harmonics are consistent with the outcomes of Eq. (5). Insets show the intensity of even-order harmonics as functions of the parameters controlling the $\hat{\mathcal{C}}_{2}$-symmetry breaking: the trigonal warping $\lambda$ for (c), and the charge imbalance $\Delta\varepsilon$ for (f).
@@ Acknowledgements
F.F. and F.G. acknowledge support from the Italian Ministry of University and Research (MUR), under Grant PRIN No. 2020JZ5N9M (CONQUEST) and Grant PRIN No. 2022TWZ9NR (STIMO). M.C. acknowledges partial support from the EU Horizon 2020 research and innovation program under Grant Agreement No. 964398 (SUPERGATE) and, together with F.F., from PNRR MUR project PE0000023-NQSTI. C.O. acknowledges support by the Italian Ministry of Foreign Affairs and International Cooperation, Grant No. PGR12351 (UTLRAQMAT) and the PNRR MUR project PE0000023-NQSTI (TOPQIN).

### v2

@@ Abstract
Figure 1: Illustration of spin-texture patterns in momentum space for high-harmonic generation. Left: in 2D systems with spin textures invariant under twofold rotational symmetry, $\hat{\mathcal{C}}_{2}$, there is no emission of even-order harmonics. Right: for spin textures breaking $\hat{\mathcal{C}}_{2}$ symmetry, the emission of even-order harmonics is allowed.
@@
This can be compared to the corresponding transformation of the spin texture, $\boldsymbol{\sigma}_{{\bf k}_{\parallel},-{\bf k}_{\perp}}^{\parallel}=-\boldsymbol{\sigma}_{{\bf k}_{\parallel},{\bf k}_{\perp}}^{\parallel}$ and $\boldsymbol{\sigma}_{{\bf k}_{\parallel},-{\bf k}_{\perp}}^{\perp}=\boldsymbol{\sigma}_{{\bf k}_{\parallel},{\bf k}_{\perp}}^{\perp}$. From this point onward, we focus on even-order harmonics. Having identified the $\hat{\mathcal{C}}_{2}$-invariant spin texture as a sufficient condition for their suppression, we now examine scenarios where spin textures break this symmetry, enabling their emission. In this context, it is essential to distinguish between the two cases in which TR symmetry is either preserved or broken.
Figure 2: Spin textures and high-harmonic response for different physical configurations. Schematic illustration of the trigonal lattice (a) with threefold rotational symmetry and antiferromagnetic pattern (d) with broken twofold rotational symmetry. Both systems possess a vertical mirror symmetry $\mathcal{M}_{x}$ along the $x$-axis, and in both cases, the incident field (green arrow) is polarized perpendicular to this direction. Blue and orange arrows represent the expected polarizations for odd- and even-order harmonics, respectively. Panels (b) and (e) show the spin textures in momentum space corresponding to the physical configuration in (a) and (d), respectively. The color legend indicates the out-of-plane spin component $\langle\sigma_{z}\rangle$, with color blue, gray and red corresponding to $\langle\sigma_{z}\rangle=-1$, 00 and $1$, respectively. Black circles mark spots in the spin textures where $\hat{\mathcal{C}}_{2}$ symmetry, as described by Eq. (4), is broken. Panels (c) and (f) show the HHG spectra (in logarithmic scale) for the time-reversal-invariant trigonal configuration and the antiferromagnetic pattern, respectively. The Hamiltonian parameters, in units of the hopping energy $t$, are: $\gamma_{R}=0.1$ for both cases; $\lambda=0.1$ in (c); and $\Delta\varepsilon=0.1$, $B=0.1$ in (f). The computations were performed at temperature $T=0$ with chemical potential $\mu=1.0$. In both cases, the incident field has intensity $|{\bf A}_{0}|=1.0$ and frequency $\Omega_{in}=0.5$. Polarizations of odd- (blue dots) and even-order (orange squares, with different shades of orange) harmonics are consistent with the outcomes of Eq. (5). The insets show the intensity of even-order harmonics (in linear scale) as functions of the parameters controlling the $\hat{\mathcal{C}}_{2}$-symmetry breaking: the trigonal warping $\lambda$ in (c) and the charge imbalance $\Delta\epsilon$ in (f). The units and color legend of even-order harmonics in each inset are consistent with those of the corresponding panel.
@@ Acknowledgements
We acknowledge valuable discussions with Jeroen van den Brink. F.F. and F.G. acknowledge support from the Italian Ministry of University and Research (MUR) under Grant PRIN No. 2020JZ5N9M (CONQUEST) and under the National Recovery and Resilience Plan (NRRP), Call PRIN 2022, funded by the European Union - NextGenerationEU, Mission 4, Component 2, Grant No. 2022TWZ9NR (STIMO)-CUP B53D23004560006. M.C. acknowledges partial support from the EU Horizon 2020 research and innovation program under Grant Agreement No. 964398 (SUPERGATE) and, together with F.F., from NRRP MUR project PE0000023-NQSTI. C.O. acknowledges support by the Italian Ministry of Foreign Affairs and International Cooperation, Grant No. PGR12351 (UTLRAQMAT) and the NRRP MUR project PE0000023-NQSTI (TOPQIN).